b) Do the following fields satisfy all four Maxwell's equations? E(t)=E_(0)sin x sin t , B(t)=B_(0)cos x cos t

1
Expert's answer
2022-02-02T12:48:26-0500

"\nabla \vec E=\frac \rho{ \varepsilon_0},"

"\nabla \vec B=0,"

"\nabla×\vec E=-\frac{\partial \vec B}{\partial t},"

"\nabla×\vec B=\mu_0(\vec J+\varepsilon_0\frac{\partial \vec E}{\partial t})."